Use the following information to answer questions 1.2 to 1.4:

Suppose the demand for a product can be represented by Qd = 100 – 5P, while

the supply is given by the equation Qs = –12 + 3P, where price is in rand. Determine the equilibrium price and quantity


1
Expert's answer
2022-03-22T19:06:40-0400

At equilibrium, demand equals supply,

Therefore, at equilibrium

"Qd=Qs"

"100-5P=-12+3P"

"100+12 = 5P+3P"

"112=8P"

Therefore, equilibrium price

=14 Rands.

"Qd=100-5(14)\n = 30"

"Qs=-12+3(14)=30"

Therefore, the equilibrium quantity is 30.